# Short-wavelength mesophases in the ground states of core-softened particles in two-dimensions

### v2

@@ Abstract
We describe the formation of short-wavelength mesophases in a two-dimensional core-softened particle system. By proposing a series of specific ansatz for each relevant phase, we performed a variational analysis to obtain the ground-state phase diagram. Our results reveal a variety of cluster lattice phases with distinct cluster orientations, alongside traditional two-dimensional Bravais lattices such as square, triangular, oblique, and rectangular structures, as well as other non-Bravais arrangements including honeycomb and kagome phases. We characterize in detail the ground-state phase transitions and identify coexistence regions between competing phases, capturing both first-order and continuous transitions. In addition, we highlight the crucial role of the competing length scales introduced by the hard-core repulsion in shaping the rich landscape of mesophases, emphasizing the interplay between intra-cluster structure and inter-cluster organization. This study provides a systematic framework that could support future investigations of classical thermal melting behavior or quantum phase transitions in similar cluster-forming systems.
  • July 2025
@@ 1 Introduction
When a hard-core repulsion or additional sharp length scale is introduced at the short distances, the tendency to form overlapping clusters can be significantly inhibited, paving the way for the formation of new modulated states with properties between solid and liquid phases, for which are sometimes called mesophases [3]. At higher densities, such modulated phases includes long-wavelength stripes, clusters and holes [2, 13, 12, 48]. At lower densities, however, these phases are characterized by short-wavelength modulations [49, 50, 51, 52, 53], and includes phases such as square lattices [54], as well as stripes of one particle width, kagome and honeycomb phases [24, 55, 56, 57], among many others. In particular, the cluster phases with a small occupation number $n$ are often identified as dimers for $n=2$ [58], trimers ($n=3$), tetramers ($n=4$) [59], and so forth.
@@
In situations when the effective hard-core size is much lower than the spacing of the lattice formed by the arrangement of the clusters themselves, the particles belonging to a cluster develop an internal structure that minimizes the hard-core repulsion energy - typically a BCC or triangular lattice in three or two dimensions, respectively. Meanwhile, in this case, the long-distance physical properties of the cluster lattice remains essentially unaltered in comparison to the corresponding soft-core particle system [48]. This kind of behavior naturally disappears as soon as the size of the clusters becomes comparable to the distance between them. In such regime the intra and inter cluster interactions are not well separated by an energy scale and, consequently, the behavior of the system becomes much richer, which can be explored already at the ground state level [63]. This scenario is exactly the one explored in Refs. [60] and [64], where the authors investigated the properties of a self-assembled crystalline dimer phase in a regime where the hard-core size of the particles is comparable to the distance between dimers forming the lattice. Interestingly, the authors shows that the classical ground state is given by a “nematic” state in which dimers are oriented in a non-trivial direction with respect to the dimers lattice.
@@
Within this context, our work aims to present a systematic study of the self-assembling process in a 2D soft-core cluster forming model frustrated by a hard-core interaction. In order to do so, we consider a two dimensional GEM-4 model combined with an inverse power-law potential, as an effective hard-core interaction, and compute the exact classical ground state phase diagram, using as running parameters the density and the relative intensity between the GEM-4 interaction and the short-range power-law repulsion. Our results shows that the soft-core scenario, when the transition between clusters with increasing occupancy occurs as the density is increased, is significantly modified even when the size of the hard-core is relatively small. To be specific, the maximum cluster occupancy attained decreases rapidly with the size of the hard-core while a plethora of new single particle phases, not present in the phase diagram of the GEM-4 model, are developed. Finally, we would like to remark that, differently from previously studied scenarios in which the interparticle distance is much smaller than the distance between clusters, in our case the orientational order of the clusters is a non-trivial feature of the self assembling process. This makes the orientational order of the cluster states to vary along the phase diagram.
The paper is organized as follows, in Sect. 2 we introduce the model studied and the method employed for the determination of the classical ground state of the system. The Sect. 3 shows the results obtained for the phase diagram of the system, initially implementing the direct energy minimization for pure phases only and, in a second moment, including the possibility of coexistence between different phases. Moreover, a detailed description of the phases stabilized by the system is presented making evident the rich physical behavior of the system studied. Finally, Sect. 4 is devoted to present our final remarks and conclusions.
@@ 2.1 Core-softened repulsive potential model
As mentioned in the previous section, we consider a two-dimensional system of classical interacting particles with a core-softened repulsive pair potential of the form:
@@
with $\alpha=4$ and $\gamma=6$. The first term in (1) represents the soft part of the pairwise potential, known as the generalized exponential model (GEM-$\alpha$), which is a bounded potential allowing for the formation of cluster-crystal ground states, where clusters are arranged on a triangular lattice and each cluster is formed by overlapping particles, with the occupation number increasing with density for $\alpha>2$ [65, 45, 47]. This ultra-soft repulsion is combined with an inverse-power-law potential, the second term in (1), which effectively acts as a hard-core repulsion, preventing particles from overlapping and thereby enabling the emergence of a variety of fine-structured ground states, partially characterized in this work.
The parameter $C$ controls the relative strength of the hard-core repulsion with respect to the ultra-soft interaction, thereby regulating the different spatial scales on which these two interactions act. For very small $C$ (in the limit $C\to 0$), the hard-core repulsion acts on very short length scales, and the ground state of overlapping particle clusters is replaced by one in which the clusters acquire a small size, composed of particles arranged in a triangular lattice, with the distance between clusters being much larger than the distance between particles within each cluster [48]. In the opposite limit, when $C\gtrsim 0.16$, the ultra-soft repulsion is dominated by the hard-core potential and the ground-state becomes the familiar triangular lattice.
@@ 2.2 From generic to specific ansatz
In order to identify candidates for possible short-wavelength ground-states we have performed molecular dynamics simulations. These were carried out in a NVT ensemble through an overdamped Langevin equation in dimensionless form, given by:
$\dot{\textbf{r}}_{i}=-\sum_{j}\nabla_{i}V($ $\left|\textbf{r}_{i}-\textbf{r}_{j}\right|)+\sqrt{2T}\,\xi_{i}(t),$ (2)
where $T$ is the temperature of the thermal bath and $\xi_{i}(t)$ is a white noise with $\left<\xi_{i}(t)\right>=0$ and $\left<\xi_{i}(t)\xi_{j}(t^{\prime})\right>=\delta_{ij}\delta(t-t^{\prime})$. These simulations were performed using 400 particles, whose equations of motion were integrated using a stochastic RK4 algorithm [66], with the addition of the parallel tempering technique [67, 68] in order to overcome metastable states. These exploratory results are not presented here and have not limited in any aspect our results, but were instrumental to initiate the construction of the ground-state ansatz for each phase.
The simulations suggested that, in the low-density region of interest $\rho<2.5$, corresponding to short-wavelength ground states, the cluster phases have occupation numbers between 2 and 4. To explore as many possible cluster ground states as feasible with energy minimization, we began by considering a generic ansatz, illustrated for cluster 2 in Figure 1. Rather than assuming a triangular lattice of clusters, we allowed a general angle $\theta$ between primitive translation vectors and variable lattice parameters along each primitive direction, described by aa along the x-axis and $b$ along the $\theta$ direction. The distance between particles within a cluster is denoted $d_{1}$, oriented along angle $\phi_{1}$. Additionally, we allowed for a sublattice displaced from the first lattice by a distance $b_{2}$ at angle $\theta_{2}$, with its own cluster size and orientation described by $d_{2}$ and $\phi_{2}$. Similar general ansatz were constructed for clusters with occupation numbers 3 and 4.
@@
$\frac{E}{N}=\frac{1}{2ns}\left(\sum_{l,m}^{\infty}\sum_{q,q^{\prime}}^{n}\sum_{r,r^{\prime}}^{s}V(|\textbf{R}_{lm,r}+\textbf{u}_{q,r}-\textbf{u}_{q^{\prime},r^{\prime}}|)\right),$ (3)
where $n$ is the cluster size and $s$ is the number of sub-lattices (in our case $s=2$). Moreover, the set $\{\textbf{R}_{lm,r}\}$ and $\{\textbf{u}_{q,r}\}$ represent the positions of the center of the clusters and the positions of the particles in the cluster respect to its center for each sub-lattice $r$, respectively. The indexes $l$ and $m$ serve to identify a specific lattice vector once we specify the corresponding lattice vector basis. As an example, to reproduce the configuration in Figure 1, we should set:
$\textbf{R}_{lm,1}$ $\equiv l\textbf{a}+m\textbf{b},$ (4)
$\textbf{R}_{lm,2}$ $\equiv l\textbf{a}+m\textbf{b}+\textbf{b}_{2},$
where the set $\{\textbf{a},\textbf{b}\}$ and $\textbf{b}_{2}$ represent the oblique vector basis of the lattice and relative position of the second sublattice respectively (see Fig. 1).
@@
$\frac{E}{N}=\frac{1}{2n}\left(\sum_{l,m}^{\infty}\sum_{q,q^{\prime}}^{n}V(|\textbf{R}_{lm}+\textbf{u}_{q}-\textbf{u}_{q^{\prime}}|)\right),$ (6)
where $\textbf{R}_{lm}\equiv l\textbf{a}+m\textbf{b}$ and $\textbf{u}_{q+1}=\{d\cos(\phi+q\pi/3),d\sin(\phi+q\pi/3)\}$, for $q\geq 0$ and $\textbf{u}_{0}=\{0,0\}$. Here the index convention is the same as in Eq.(3). Note that in general, especially for larger clusters, it may be necessary to modify the $\textbf{u}_{q}$, however, up to cluster 4, our parametrization will accurately describe cluster geometries.
The only exception to the description given by eq. (6) is the Cluster 3a ansatz. In this case clusters alternate orientation along the b direction introducing a secondary sub-lattice in that direction. Nevertheless, in this case it is simple to realize that the total interaction energy of a trimer in both sub-lattices are equal. This results is straightforward if we observe that the trimers in both sub-latices are equivalent by a lattice reflection using a as symmetry axis. For this reason, to evaluate the total energy we simply consider the value obtained for a single type of trimer.
@@
Taking into account this kind of constraint, we have as independent variational parameters $a$, $\theta$ and $d$ for all cluster phases, $a$ and $\theta$ for the oblique phase and just $a$ for the rectangular phase. On the other hand, the triangular, square, honeycomb and kagome phases are fully determined by the density, without any minimization. In order to enforce each ansatz to represent a single type of configuration we should impose additional constraints to the variational parameters as presented in 2.

### v3

@@ Abstract
We describe the formation of short-wavelength mesophases in a two-dimensional core-softened particle system. By proposing a series of specific ansatz for each relevant phase, we performed a variational analysis to obtain the ground-state phase diagram. Our results reveal a variety of cluster lattice phases with distinct cluster orientations, alongside traditional two-dimensional Bravais lattices such as square, triangular, oblique, and rectangular structures, as well as other non-Bravais arrangements including honeycomb and kagome phases. We characterize in detail the ground-state phase transitions and identify coexistence regions between competing phases, capturing both first-order and continuous transitions. In addition, we highlight the crucial role of the competing length scales introduced by the hard-core repulsion in shaping the rich landscape of mesophases, emphasizing the interplay between intra-cluster structure and inter-cluster organization. Finally, our analytical results are confronted with extensive molecular dynamics simulations, which interestingly show the existence of decagonal and dodecagonal quasicrystalline phases in regions of the phase diagram that exhibit a high degree of frustration. This study provides a systematic framework that could support future investigations of classical thermal melting behavior or quantum phase transitions in similar cluster-forming systems.
  • December 2025
@@ 1 Introduction
When a hard-core repulsion or additional sharp length scale is introduced at the short distances, the tendency to form overlapping clusters can be significantly inhibited, paving the way for the formation of new modulated states with properties between solid and liquid phases, which are sometimes called mesophases [3]. At higher densities, such modulated phases includes long-wavelength stripes, clusters and holes [2, 13, 12, 48]. At lower densities, however, these phases are characterized by short-wavelength modulations [49, 50, 51, 52, 53], and includes phases such as square lattices [54], as well as stripes of one particle width, kagome and honeycomb phases [24, 55, 56, 57], among many others. In particular, the cluster phases with a small occupation number $n$ are often identified as dimers for $n=2$ [58], trimers ($n=3$), tetramers ($n=4$) [59], and so forth. More exotic phases can be found in such systems, as is the case of sigma phases [26, 24], considered as periodic approximant for quasicrystalline phases, as well as quasicrystalline phases with $Q$-fold rotational symmetry, with the more common $Q=12$ corresponding to dodecagonal quasicrystal [26, 24, 25, 60].
@@
In situations when the effective hard-core size is much lower than the spacing of the lattice formed by the arrangement of the clusters themselves, the particles belonging to a cluster develop an internal structure that minimizes the hard-core repulsion energy - typically a BCC or FCC lattice in three dimensions [47, 64] or a triangular lattice in two dimensions. Meanwhile, in this case, the long-distance physical properties of the cluster lattice remains essentially unaltered in comparison to the corresponding soft-core particle system [48]. This kind of behavior naturally disappears as soon as the size of the clusters becomes comparable to the distance between them. In such regime the intra and inter cluster interactions are not well separated by an energy scale and, consequently, the behavior of the system becomes much richer, which can be explored already at the ground state level [65]. This scenario is exactly the one explored in Refs. [61] and [66], where the authors investigated the properties of a self-assembled crystalline dimer phase in a regime where the hard-core size of the particles is comparable to the distance between dimers forming the lattice. Interestingly, the authors show that the classical ground state is given by a “nematic” state in which dimers are oriented in a non-trivial direction with respect to the dimers lattice.
@@
Within this context, our work aims to present a systematic study of the self-assembling process in a 2D soft-core cluster forming model frustrated by a hard-core interaction. In order to do so, we consider a two-dimensional model with a core-softened potential that combines the generalized exponential model with exponent 4 (GEM-4) and an inverse power-law potential with exponent 6 – a similar core-softened potential was considered in a previous study [48] that investigated cluster-crystal ground-states in much higher densities and much lower effective hard-core sizes than those considered in the present work. We compute in detail the classical ground-state phase diagram, using density and the relative strength of the two repulsive terms as running parameters. Our results show that the hard-core interaction significantly modifies the soft-core scenario of clusters with increasing occupancy. The maximum cluster occupancy decreases rapidly with the hard-core size, while a plethora of new single-particle phases emerge, which are not present in the pure GEM-4 model. Furthermore, we explore the intermediate regime where both interactions act on comparable scales, making the orientational order of the clusters a non-trivial feature that varies along the phase diagram.
The paper is organized as follows. In Sect. 2 we introduce the model and the methods used to determine the classical ground state. Sect. 3 presents a detailed description of the stabilized phases and their location in the phase diagram, first from direct energy minimization of pure phases and then allowing for phase coexistence; this section also includes a correction to the phase diagram based on simulations. Finally, Sect. 4 is devoted to present our final remarks and conclusions.
@@ 2.1 Core-softened repulsive potential model
We consider a two-dimensional system of classical particles interacting via a pair potential composed of two competing repulsive terms: an ultrasoft and bounded interaction that favors particle clustering, and a short-range interaction that penalizes particle overlap. It can be generally expressed as:
$U(r)=\epsilon_{s}\exp\left(-(r/R)^{\alpha}\right)+\epsilon_{s}(r_{0}/r)^{\gamma},$ (1)
where $\epsilon_{s}$ controls the energy scale, and the first term represents the soft repulsion with typical length scale $R$, while the second term corresponds to the power-law hard-core repulsion with typical length scale $r_{0}$, characterizing the effective particle size. By expressing energies in units of $\epsilon_{s}$ and distances in units of $R$, we rewrite the potential in the following dimensionless form, which is used throughout this work:
@@
with the particular values $\alpha=4$ and $\gamma=6$. The dimensionless parameter $C=(r_{0}/R)^{\gamma}$ controls the relative strength of the hard-core repulsion. In our analysis and phase diagrams, we often use the parameter $\ell_{C}\equiv-\log C$, as it conveniently represents the relative hard-core strength across different orders of magnitude. The first term in (2) represents the soft part of the pairwise potential, known as the generalized exponential model (GEM-$\alpha$), which is a bounded potential allowing for the formation of cluster-crystal ground states, where clusters are arranged on a triangular lattice and each cluster is formed by overlapping particles, with the occupation number increasing with density for $\alpha>2$ [64, 45, 47]. This ultra-soft repulsion is combined with an inverse-power-law potential, the second term in (2), which effectively acts as a hard-core repulsion, preventing particles from overlapping and thereby enabling the emergence of a variety of fine-structured ground states, partially characterized in this work.
Since the parameter $C$ controls the relative strength of the hard-core repulsion with respect to the ultra-soft interaction, it thereby regulates the different spatial scales on which these two interactions act. For very small $C$ (in the limit $C\to 0$), the hard-core repulsion acts on very short length scales, and the ground state of overlapping particle clusters is replaced by one in which the clusters acquire a small size, composed of particles arranged in a triangular lattice, with the distance between clusters being much larger than the distance between particles within each cluster [48]. In the opposite limit, for $C\gtrsim 0.16$ ($\ell_{C}\lesssim 0.80$, as shown by our results in the following section in Fig. 4), the ultra-soft repulsion is completely dominated by the hard-core potential and the ground-state becomes the familiar triangular lattice.
@@
2.2 Molecular dynamics simulations
In order to identify candidates for possible short-wavelength ground-states, we performed extensive molecular dynamics simulations. To efficiently explore the configuration space and overcome metastable states, we employed a combination of parallel tempering and simulated annealing techniques. These exploratory simulations were instrumental in identifying the relevant phases and constructing the specific ansatz for our variational analysis.
The simulations were performed in a NVT or canonical ensemble, where we control the number of particles, the volume (area) and temperature of the system. The contact with a thermal reservoir in the dynamics of the particles is accomplished through the following Langevin equation in dimensionless form:
$\mathbf{\ddot{r}}_{i}=-\sum_{j\neq i}\mathbf{\nabla}_{i}V(\mathbf{r}_{i}-\mathbf{r}_{j})-\gamma\mathbf{\dot{r}}_{i}+\sqrt{2\gamma T}\,\mathbf{f}_{i}(t),$ (3)
where each Cartesian component of $\mathbf{f}_{i}(t)$ corresponds to a white noise with zero mean and unit variance, $\gamma$ denotes the dimensionless viscous damping of the thermal bath and $T$ corresponds to the dimensionless temperature (measured in units of $\epsilon_{s}/k_{B}$). The equations of motion were integrated using a Verlet-type stochastic scheme, specifically the G–JF algorithm [67].
Simulations of the system examined in this work exhibit a tendency to become trapped in metastable states due to the inherent frustration arising from the competing length scales in the pair potential. The degeneracy of the ground states combined with finite-size effects further slow convergence to equilibrium, by promoting the formation of arrested phase boundaries and defect structures. To mitigate these issues, we employ a parallel tempering technique (also known as the replica-exchange algorithm), which efficiently explores the configuration space near phase transitions [68, 69] and facilitates escape from metastable states at low temperatures by allowing replicas to access higher thermal energies. These simulations are then followed by an annealing of the replica associated with the lowest temperature down to temperatures approaching zero. This hybrid approach – together with the use of full inertial Langevin dynamics – has proven effective in equilibrating low-temperature phases. While parallel tempering alleviates major equilibration barriers, the final annealing step suppresses thermal fluctuations, thereby revealing the underlying ground-state pattern.
[…]
Figure 1: Cluster crystal low temperature molecular dynamics configurations corresponding to dimer (a,b) and trimer (c, d) phases, for systems with $N=900$ and $N=1200$ particles, respectively. The corresponding values of the density, relative strength of the hard-core interaction and temperature are provided above each configuration. The associated structure factors of the configurations are shown in the insets.
We have performed simulations with a number of particles ranging $N$ from 400 to 1200 particles using $m=1$, $\gamma=0.05$ and $dt=0.005$ in eq. (3). For the parallel tempering we have used approximately $n_{T}\simeq 100$ temperatures, distributed according to the Hukushima-Nemoto algorithm [68] such that the swapping ratio between neighboring replicas is almost constant and independent of the temperature values – the highest temperature values ranging from $0.06$ to $0.12$, depending on the phase diagram region. Figure 1 shows representative configurations obtained from these simulations in the case of cluster crystal phases with occupation number 2 and 3. As an initial survey, we performed a broad sweep of the phase diagram using increments $\Delta\rho=0.05$ and $\Delta\ell_{C}=0.05$, enabling us to identify the overall structure of the phase diagram with moderate resolution. These final configurations from the simulations served as an inspiration for the construction of the several variational ansatz developed to investigate the ground-state phase diagram in this system.
Nevertheless, as will become clear shortly, more general ansatz than those observed in the simulations were constructed, so the simulations have not limited our results in any aspect. While one could attempt to sketch a phase diagram based solely on numerical simulations, this strategy has notable drawbacks: as discussed earlier, simulations may settle into metastable configurations, which makes it difficult to obtain a reliable comparison between the energies of different ground-states. In addition, simulation results alone do not produce a reliable characterization of the nature of phase transitions and identification of possible coexistence regions — features that become accessible when the energy of each ground-state candidate is analytically determined, as it will become clear throughout the present work in the following sections. Conversely, the simulations are particularly valuable for exposing quasiperiodic phases, for which constructing a suitable ansatz is far from trivial. The quasicrystalline phases obtained from simulations will be examined in more detail at the end of Sect. 3.
@@ 2.3 From generic to specific ansatz
The simulations results suggested that, in the low-density region of interest $\rho<2.5$, corresponding to short-wavelength ground states, the cluster phases have occupation numbers between 2 and 4. To explore as many possible cluster ground states as feasible with energy minimization, we began by considering a generic ansatz, illustrated for cluster 2 in Figure 2. Rather than assuming a triangular lattice of clusters, we allowed a general angle $\theta$ between primitive translation vectors and variable lattice parameters along each primitive direction, described by $a$ along the x-axis and $b$ along the $\theta$ direction. The distance between particles within a cluster is denoted $d_{1}$, oriented along angle $\phi_{1}$. Additionally, we allowed the existence of a second sub-lattice, identical to the original one, but displaced by a distance $b_{2}$ at angle $\theta_{2}$, with its own cluster size and orientation described by $d_{2}$ and $\phi_{2}$. Similar general ansatz were constructed for clusters with occupation numbers 3 and 4.
@@
$\frac{E}{N}=\frac{1}{2ns}\left(\sum_{l,m}^{\infty}\sum_{q,q^{\prime}}^{n}\sum_{r,r^{\prime}}^{s}V(|\textbf{R}_{lm}+\textbf{s}_{r}-\textbf{s}_{r^{\prime}}+\textbf{u}_{q,r}-\textbf{u}_{q^{\prime},r^{\prime}}|)\right),$ (4)
where $n$ is the cluster size and $s$ is the number of sub-lattices (in our case $s=2$). Moreover, the set $\{\textbf{R}_{lm}\}$ and $\{\textbf{u}_{q,r}\}$ represent the positions of the center of the clusters in a sub-lattice, while $\textbf{s}_{r}$ represent the position of the corresponding sub-lattice respect to the first sub-lattice (so $\textbf{s}_{1}=\textbf{0}$). Naturally, the indexes $l$ and $m$ serve to identify a specific lattice vector once we specify the corresponding lattice vector basis. As an example, to reproduce the configuration in Figure 2, we should set:
$\textbf{R}_{lm}$ $\equiv l\textbf{a}+m\textbf{b},$ (5)
$\textbf{s}_{2}$ $\equiv\textbf{b}_{2},$
where the set $\{\textbf{a},\textbf{b}\}$ and $\textbf{b}_{2}$ represent the oblique vector basis of the lattice and the relative position of the second sub-lattice, respectively (see Fig. 2).
@@
$\frac{E}{N}=\frac{1}{2n}\left(\sum_{l,m}^{\infty}\sum_{q,q^{\prime}}^{n}V(|\textbf{R}_{lm}+\textbf{u}_{qlm}-\textbf{u}_{q^{\prime}00}|)\right),$ (7)
where $\textbf{R}_{lm}\equiv l\textbf{a}+m\textbf{b}$ and $\textbf{u}_{(q+1)lm}=\{d\cos(\phi_{lm}+q\pi/3),d\sin(\phi_{lm}+q\pi/3)\}$, for $q\geq 0$. The site dependent phases $\phi_{lm}$ are taken taken as a single variational parameter for all Cluster 3 phases except for the case of Cluster 3a phase in which we have two different equivalent sub-lattices with opposite cluster orientation. Here the index convention is the same as in Eq.(4). Note that in general, especially for larger clusters, it may be necessary to modify the $\textbf{u}_{q}$, however, up to cluster 4, our parametrization will accurately describe cluster geometries. Despite the fact that Cluster 3a phase have two different sub-lattices the validity of Eq. (7) in this case stems from the fact that both sub-lattices forming the crystal are equivalent by a lattice reflection using a as symmetry axis.
@@
Taking into account this kind of constraint, we have as independent variational parameters $a$, $\theta$ and $d$ for all cluster phases, $a$ and $\theta$ for the oblique phase and just $a$ for the rectangular phase. On the other hand, the triangular, square, honeycomb and kagome phases are fully determined by the density, without any minimization. In order to enforce each ansatz to represent a single type of configuration we should impose additional constraints to the variational parameters as presented in Figure 3. For each ansatz, we minimized its energy with respect to its variational parameters at each point in the ($\rho,\ell_{C}$) plane. To efficiently construct the phase diagram, for each $\ell_{C}$ we implemented a sweeping protocol in density, where the optimized parameters for each ansatz at a given $\rho$ served as the initial condition for the same ansatz in the next point $\rho+\Delta\rho$.